### v1

# Specialist vision-language models for clinical ophthalmology

Abstract
Clinicians spend a significant amount of time reviewing medical images and transcribing their findings regarding patient diagnosis, referral and treatment in text form. Vision-language models (VLMs), which automatically interpret images and summarize their findings as text, have enormous potential to alleviate clinical workloads and increase patient access to high-quality medical care. While foundational models have stirred considerable interest in the medical community, it is unclear whether their general capabilities translate to real-world clinical utility. In this work, we show that foundation VLMs markedly underperform compared to practicing ophthalmologists on specialist tasks crucial to the care of patients with age-related macular degeneration (AMD). To address this, we initially identified the essential capabilities required for image-based clinical decision-making, and then developed a curriculum to selectively train VLMs in these skills. The resulting model, RetinaVLM, can be instructed to write reports that significantly outperform those written by leading foundation medical VLMs in disease staging (F1 score of 0.63 vs. 0.11) and patient referral (0.67 vs. 0.39), and approaches the diagnostic performance of junior ophthalmologists (who achieve 0.77 and 0.78 on the respective tasks). Furthermore, in a reader study involving two senior ophthalmologists with up to 32 years of experience, RetinaVLM’s reports were found to be similarly correct (78.6% vs. 82.1%) and complete (both 78.6%) as reports written by junior ophthalmologists with up to 10 years of experience. These results demonstrate that our curriculum-based approach provides a blueprint for specializing generalist foundation medical VLMs to handle real-world clinical tasks.
footnotetext: denotes shared senior authorship
1 Introduction
Figure 1: We introduce RetinaVLM, the first specialist medical generative vision-language model (VLM). (a) Using a needs-driven training curriculum, we trained RetinaVLM in specialist medical skills that medical foundation VLMs are currently lacking (b) RetinaVLM is able to process retinal optical retinal optical coherence tomography (OCT) images and flexibly respond to text-based queries. (c) Its abilities entail the analysis of imaging biomarkers of age-related macular degeneration (AMD), disease staging, and the referral for treatment.
Medical images are central to many clinical decisions regarding patient diagnosis, referral, and treatment. Clinicians spend a significant amount of time transcribing image-based decisions into text in order to store and communicate their findings [1, 2]. Visual-language models (VLM), which automatically interpret medical images and generate detailed textual descriptions, have enormous potential to alleviate clinical workloads and increase patient access to high-quality medical care [3, 4]. To date, the majority of medical VLMs have been trained to output a finite set of pre-determined textual responses [5, 6, 7, 8]. Only recently, the combination of large language models (LLM) with medical vision encoders has led to the development of more powerful and versatile generative VLMs that are able to write comprehensive text reports or answer complex questions [9, 10, 11].
This current generation of medical language models is fueled by vast amounts of unstructured training data that is extracted from medical textbooks, scientific publications or social media posts of healthcare professionals [9, 6, 10]. These foundation language models have stirred considerable interest among the medical community for their expert-level performance on standardized medical question-answering tasks, such as licensing exams and case studies [12, 13]. However, it is unclear whether this general performance translates to clinical utility in specialist medical domains [14]. Despite its impressive scale, the training data of foundation language models has been collected agnostically towards their downstream application and inherently lacks specialist information related to the challenges faced by clinicians in their daily practice.
In this study, we identify this missing piece in foundation models with the aim of developing generative medical VLMs with real-world clinical utility. We propose to deconstruct clinical problems into sets of mandatory capabilities required for their resolution and selectively train VLMs in these skills. We demonstrate the feasibility of this approach in ophthalmology. To this end we introduce RetinaVLM, the first specialist medical generative VLM (see Figure 1a). RetinaVLM is trained using a dedicated curriculum that is specific to the clinical management of age-related macular degeneration (AMD), the leading cause of blindness in the elderly [15, 16]. It is able to process optical coherence tomography (OCT) images of the retina and flexibly respond to instructions and questions (see Figure 1b). In extensive experiments, we evaluate RetinaVLM’s utility and versatility regarding disease staging, patient referral and biomarker analysis in AMD (see Figure 1c).
2 Results
2.1 RetinaVLM, a specialist vision-language model for retinal image analysis
RetinaVLM combines two main components: an ophthalmic vision encoder that processes input OCT images, and a generative LLM that handles textual instructions and outputs the corresponding responses (see Figure 1a). The vision encoder is based on our previous work, in which we found it to perform on par with RETFound [17], a large foundation model for retinal image analysis [18]. We use Meta’s Llama 3 as generative LLM, the best performing model that is openly available at the time of this study [19]. However, without additional training it lacks specialist knowledge related to the analysis of OCT images and the clinical management of AMD. Both these deep neural networks have already been pre-trained on large OCT and natural language datasets, respectively, and we only finetune them in the scope of this study. Additional details about the network architecture and training are included in Section 5.2.
2.2 A curriculum to encode specialist ophthalmological knowledge in vision-language models
An intuitive strategy to specialize VLMs while preserving their ability to flexibly interact with text queries is to provide them with a set of medical images and corresponding question-answer pairs. They are then optimized based on the similarity of their predicted answers to the ground truth. However, visual question-answering (VQA) datasets do not exist for most medical specializations, including ophthalmology.
Together with a large team of ophthalmologists, which are involved with the patient care and academic research of AMD, we defined a set of required capabilities for accurate image-based clinical management of AMD. They include the identification of AMD biomarkers in OCT images, the linking of these to the AMD disease stage, and ultimately deciding on the required referral and treatment of the patient. We then curated a training curriculum, which consists of 41,926 OCT images, and 479,710 visual questions and answers to progressively specialize VLMs in these capabilities.
Figure 2: We curated a two-part curriculum to specialize medical VLMs for clinical use. (a and b) Based on a retrospectively collected OCT imaging dataset, we created a large number of tabular reports as well as a small number of comprehensive textual reports. (c and d) We then used an independent LLM to automatically generate visual question-answers (VQA) based on these reports. (e and f) This yielded two VQA datasets, the first on basic imaging biomarkers of AMD and the second covering more advanced clinical skills. (g and h) Finally, we trained two specialist medical generative VLMs, RetinaVLM-Base and RetinaVLM-Specialist, using either the first or both VQA datasets.
Curriculum part 1: Introduction to retina
The first part of the curriculum, named Introduction to retina, primarily covers the appearance of the retina and AMD biomarkers in OCT images. Using automated data collection, we obtained tabular reports for 41,926 retrospectively collected OCT images of AMD patients (see Figure 2a). Each report describes the visible biomarkers, patient’s diagnosis, visual acuity and demographic information in 34 data fields. A full description of the OCT dataset can be found in Section 5.1.1, the list of all tabular data fields and example tabular reports in Figure 7, and the methodology for their automated procurement in Section 5.4.1.
Next, we tasked an independent LLM to generate question-answer pairs based on these reports (see Figure 2c). The model processed the content of the tabular reports – but not the OCT images – to output a numbered list of question-answer pairs. We generated an average of ten question-answer pairs per report that are mostly related to the presence or absence of specific biomarkers (see Figure 2e). The LLM was instructed to create both closed-ended ’yes or no’ style questions, and simple open-ended questions. Detailed information on the LLM setup can be found in Section 5.4.1.
This automated approach allowed us to generate a large dataset of 408,545 question-answer pairs. However, the questions were limited in scope to the set of biomarkers documented by the tabular reports. Training on these yielded the first of two specialist VLMs, RetinaVLM-Base (see Figure 2g).
Curriculum part 2: Advanced retinal specialism
The second part of the curriculum, named Advanced retinal specialism, builds on top of the first part to link imaging biomarkers to AMD stage and the recommended course of treatment. As this reasoning cannot be fully conveyed via tabular information, we tasked two ophthalmologists with 3 and 10 years of experience, respectively, to create comprehensive textual reports for a subset of 330 OCT images (see Figure 2b). The ophthalmologists were asked to primarily describe the main pathological biomarkers related to AMD while also noting any other observations regarding the retinal anatomy. This task yielded high-quality reports that go beyond the short notes that are typically written by ophthalmologists in their clinical routine. Instructions given to the ophthalmologists as well as a set of sample reports are provided in Section 5.4.2 and Figure 8, respectively.
Similar to before, an independent LLM was then employed to automatically generate question-answer pairs based on the reports (see Figure 2d). Due to the substantially increased depth and scope of the full-text reports compared to the tabular ones, we used several advanced LLM instructions to create 216 diverse question-answer pairs per image on average (see Figure 2f). These cover additional biomarkers and sub-categorize them based on their size, type, and location. Other question-answer pairs are related to the causal relationship between biomarkers and six AMD disease stages as well as three levels of patient referral urgency. Moreover, the question-answer pairs were more varied in their structure in order to preserve interactive capabilities of the foundation LLM. For example, some queries asked to summarize the existing reports or provide several answers in succession. An example interaction with the LLM to generate question-answer pairs with the LLM is shown in Figure 9. Furthermore, a list of all the LLM instructions is provided in Section A.2 and example question-answers yielded by this approach are shown in the ‘part 2’ section of Figure 10.
This resulted in a dataset of 71,165 advanced question-answer pairs. By further training RetinaVLM-Base on the second part of the curriculum, we obtained our most performant VLM for the clinical management of AMD, RetinaVLM-Specialist (see Figure 2h).
2.3 RetinaVLM-Specialist outperforms foundation models and approaches junior ophthalmologists in AMD disease staging and report writing
Figure 3: (a) Comparison of the ability of four VLMs to write reports on retinal OCT images and derive the AMD stage. (b) Overall staging accuracy for each model was calculated using micro F1 scores with 95% CI, with tests of statistical significance calculated using McNemar’s test. (c) Confusion matrices between the senior ophthalmologists’ assessments (rows) against the image-based clinical decision maker’s prediction (columns). (d) Qualitative comparison of reports written by human ophthalmologists and RetinaVLM-Specialist with text markings highlighting findings regarding biomarker observations and disease stage.
Estimating the disease stage is crucial to patient management as it allows clinicians to monitor and treat patients using standardized protocols. We assessed the ability of four different generative VLMs to determine the AMD disease stage from retinal OCT images. Specifically, we compared two medical foundation VLMs, Med-Flamingo [10] and LLaVA-Med [9], to our two specialist VLMs, RetinaVLM-Base and RetinaVLM-Specialist. Using a testing dataset of previously unseen 276 OCT images, VLMs were tasked to write reports that describe the OCT image before classifying the patient into one of six disease stages (see Figure 3a). The model predictions were compared to ground truth labels obtained from ophthalmologists. Each image was initially graded by two out of six junior ophthalmologists, whose experience in the field ranges from 2 to 15 years. Inter-rater disagreements were resolved by a panel of two senior ophthalmologists with 25 and 32 years of experience, respectively. For additional methodological details, including the instruction given to the VLMs to generate these reports, see Section 5.8.1.
We found that our intermediate RetinaVLM-Base model already performs significantly better than both foundation VLMs, which lack the ophthalmological specialism to stage disease (see Figure 3b). The most performant foundation VLM, Med-Flamingo, achieved a F1 score of 0.11. This was markedly outperformed by our advanced RetinaVLM-Specialist model, scoring at 0.63. This approached, but did not match, the accuracy of the junior ophthalmologists who achieved an F1 score of 0.78. We analyze this discrepancy in detail in Section 3. Both foundation VLMs and RetinaVLM-Base returned a substantial number of invalid reports that did not conclude with one of the six disease stages (see Figure 3c). Conversely, all generated reports by RetinaVLM-Specialist were valid. Similar to human experts, RetinaVLM-Specialist struggled the most when diagnosing wet inactive AMD. We attribute this to the high number of shared imaging biomarkers that indicate either intermediate and late-wet forms of AMD, which sometimes leads to misdiagnosis by both ophthalmologists and RetinaVLM-Specialist (see Figure 3d). This was despite the identification of emerging features related to inactive late wet AMD (a small amount of hyperreflective material, or scar tissue) by RetinaVLM. Four more examples of success and failure cases of RetinaVLM-Specialist are shown in Figure 11a. Moreover, full numerical results as well as the confusion matrix for Med-Flamingo are shown in Figure 12 and 13, respectively.
Figure 4: (a) Correctness, completeness and conciseness of the reports written by LLaVA-Med, RetinaVLM-Specialist and junior ophthalmologists. Reports were scored for on each of the three criteria by senior ophthalmologists using a five-point Likert scale. (b) Representative reports with ratings by one of the senior ophthalmologists.
Eighty-four of the generated reports were scored by the two senior ophthalmologists for their correctness, completeness, and conciseness. They were shown 28 reports written by LLaVA-Med, 28 by RetinaVLM-Specialist, and 28 by the two annotating junior ophthalmologists in random order without knowledge of the author. For each report, the senior ophthalmologist first reviewed the corresponding OCT image before rating the generated report in the three criteria on a five point Likert scale [20].
Echoing our previous findings, the senior ophthalmologists observed that LLaVA-Med failed to comprise factually correct image reports (see Figure 4a), even though it uses specialist terminology that may give the reports the initial appearance of being composed by an ophthalmologist (see Figure 4b). LLaVA-Med also failed to keep reports brief, despite being explicitly instructed to do so.
Conversely, the reports generated by RetinaVLM-Specialist were mostly graded as correct and complete and of similar quality as those written by junior ophthalmologists. Specifically, out of the 28 reports, the senior ophthalmologists either agreed or strongly agreed that 22 (78.6%) written by RetinaVLM-Specialist, and 23 (82.1%) written by junior ophthalmologists, were correct in their observations and conclusions. Similarly, they found that for 20 (71.4%) of the images, neither of these report writers missed any features that were relevant to AMD.
Reports by RetinaVLM-Specialist were deemed to be slightly less concise, where 18 out of 28 (64.3%) were rated as complete compared to 23 (82.1%) by the junior ophthalmologists. This discrepancy can be seen in the second sample in 4b, where RetinaVLM-Specialist correctly identified that the image showed a healthy retina, but also detects a small cyst that was not found in the image, resulting in a lower conciseness rating. Junior ophthalmologists wrote a concise report, but incorrectly associated subretinal drusenoid deposits with intermediate AMD. These shortcomings of both the junior ophthalmologists and RetinaVLM-Specialist are discussed in more detail in Section 3.
2.4 RetinaVLM-Specialist surpasses opticians and approaches junior ophthalmologists in AMD patient screening and referral
Figure 5: (a) Evaluation of the ability of four VLMs to assess the need for patient referral for treatment of wet AMD. (b) Overall referral accuracy was calculated using F1 score for urgent referral with a 95% CI. Tests of statistical significance were carried out using McNemar’s test. The performance of individual ophthalmologists is shown by two white points. (c) Confusion matrices between the senior ophthalmologists assessment (rows) against the image-based clinical decision maker’s referral assessment (columns). (d) Image reports written by the non-specialist optician who originally referred the patient, compared with reports of the same patient written by RetinaVLM-Specialist.
As the prevalence of AMD is expected to further increase in the upcoming decades [16], ocular screening programs are being introduced around the world. In the United Kingdom, some projects involve opticians and pharmacies that acquire and interpret OCT images. They may refer a patient to a specialist clinic, summarizing their findings and the estimated level of the patient’s risk in a letter. In the United Kingdom, treatment guidelines for AMD mandate that patients with signs of neovascularization are referred for immediate treatment within two weeks. However, non-specialists exhibit a tendency to over-diagnose these cases. An internal audit at Southampton Eye Unit found that 74.2% of the referrals made to the clinic do not have any form of treatable AMD. The processing and assessment of these false positives affects the clinic’s ability to care for the remaining patients with treatable forms of AMD.
We evaluated the ability of VLMs to assess the level of referral urgency from OCT image (see 5a). For each case, the VLMs were provided explicit referral guidelines, and asked to recommend which of three levels of referral urgency was most appropriate for the patient: no referral for healthy patients, to be seen within 18 weeks (routine referral) for patients that are at risk of progressing to active late wet AMD but do not require treatment yet, and referral within two weeks for patients with any signs of neovascularization that should be urgently referred for antiangiogenic treatment. Two junior ophthalmologists independently reviewed images of 95 patients that have previously been referred to the hospital for treatment of wet AMD. For each patient, they independently decided the most appropriate of the three levels of referral urgency, and disagreements were arbitrated by the two senior ophthalmologists. In line with previous audits, they found the false discovery rate for urgent referrals was 69.5%. We then calculated F1 scores for the highest risk patients in need of urgent referral between the VLM’s predictions and the ground truth. The full referral protocol and report generation instructions given to the VLMs are provided in Section 5.8.3.
We found that both medical foundation VLMs and Retina-Base perform worse than opticians regarding their ability to refer patients in need of urgent treatment (see Figure 5b). While Med-Flamingo failed to refer any of the 29 high-risk patients cases, LLaVA-Med and RetinaVLM-Base were ineffective for differentiating high-risk patients from low- to moderate-risk patients (see Figure 5c). RetinaVLM-Specialist was able to detect 23 out of the 29 high-risk cases that require immediate treatment. At the same time, RetinaVLM’s false discovery rate, defined as the ratio of the number of false positives over the number of predicted positives, of 42.5% is substantially lower than that of opticians at 69.5%. Owing to their ability to better differentiate moderate from high-risk cases, the human ophthalmologists had the lowest false discovery rate of 9.1%, although they simultaneously missed three more cases in urgent need for treatment.
In practice, referral letters should communicate the reason for referral by citing suspected abnormalities in the OCT image that can inform the ophthalmologist’s initial diagnostic plan. As in the conciseness study in Figure 4, RetinaVLM-Specialist sometimes documents the presence of small biomarkers that cannot be found in the image. More often, RetinaVLM-Specialist wrote an accurate imaging report but did not accurately follow the complex set of referral guidelines provided in the instruction. This led RetinaVLM-Specialist to incorrectly recommend that 17 of the moderate-risk patients potentially require treatment. However, this occurred less for the 25 low-risk patients, where RetinaVLM-Specialist correctly identified patients with little or no abnormalities, which are often referred to the treatment clinic for a second opinion by non-specialists (samples 1 and 3 in Figure 5d). Crucially, we find that RetinaVLM-Specialist is effective in the detection of intraretinal cysts and fluid that differentiate high-risk from moderate-risk patients (samples 4 and 5). Four more examples of success and failure cases of RetinaVLM-Specialist are shown in Figure 11b.
2.5 RetinaVLM accurately detects imaging biomarkers to make recommendations
Figure 6: (a) We test four VLMs on their ability to describe the presence of 10 important imaging biomarkers of AMD. (b) Overall detection accuracy was computed using F1 scores. (c) Detection sensitivity for each level of biomarker severity for the most important biomarkers. (d) Regions highlighted by RetinaVLM-Specialist that correspond with different biomarkers and disease stage assessments in its written report. Regions with greater than 25% and 50% importance are highlighted by yellow and red contours, respectively. Four additional examples are shown in Figure 14.
It is important that clinical decision makers can provide evidence for their recommendations. Disease staging reports and written referral recommendations commonly contain descriptions of the most salient biomarkers that were detected in the scan. We tested the ability of four VLMs to correctly identify the presence or absence of 10 different biomarkers related to AMD. To this end, all VLMs were tasked with writing reports for 396 OCT images that conclude by stating the presence or absence of the biomarker in question (see Figure 6a). The VLMs predictions were compared against the ground truth labels obtained from junior ophthalmologists. The instruction used to generate these biomarker focused reports is provided in Section 5.8.4.
We find that RetinaVLM-Specialist outperforms both LLaVA-Med and Med-Flamingo in the detection of seven out of the ten of main biomarkers related to AMD (see Figure 6b). Biomarkers that were more severe, larger, and more numerous were detected with higher accuracy by RetinaVLM-Specialist than less advanced presentations (see Figure 6c). Most of the smaller biomarkers, such as small amounts of intraretinal fluid, drusen and hyperreflective foci, which can be as small as 30 $\mu m$ in size [21], were detected with lower sensitivity. Overall, clinically important hallmarks of late AMD were detected with a very high sensitivity. Large volumes of subretinal and intraretinal fluid were detected in 80% and 78% of cases, respectively, and severe levels of hypertransmission in 84% of cases.
Finally, to visualize the functioning of RetinaVLM-Specialist, we calculated saliency maps based on Grad-CAM [22]. These saliency maps highlight the image regions deemed most important by the model when writing specific passages of the report (see Figure 6d). We refer to Figure 14 for four additional image reports with corresponding saliency maps. We qualitatively found that RetinaVLM-Specialist is influenced by different imaging biomarkers when writing different passages of the report, and in making its final recommendation. We observed the saliency maps were especially effective for highlighting hyperreflective material, RPE irregularities and hypertransmission.
3 Discussion
Main findings of the study
In this study, we have presented RetinaVLM, the first specialist generative visual language model in medicine. Given a retinal OCT image, RetinaVLM provides accurate, detailed textual responses related to disease staging, referral or biomarker identification of AMD. While large foundation deep learning models have been employed for retinal image analysis before [23, 18], our generative VLM is the first model that can flexibly process varied textual queries related to complex ophthalmological decisions and return detailed written responses. Through the use of language as primary communication medium, artificial intelligence systems are able to dynamically perform new tasks and meet the evolving requirements of image-based clinical decision makers.
In extensive experiments, RetinaVLM significantly outperformed state-of-the-art generative VLMs designed for medical use. We found that these are unable to interpret OCT images, derive the AMD disease stage and follow standard referral guidelines for AMD. Specifically, we have shown that in disease staging, RetinaVLM surpasses LLaVA-Med, the most performant open-source medical VLM, and is approaching the accuracy of junior ophthalmologists. When testing the ability of VLMs to screen for high-risk patients, LLaVA-Med substantially underperforms compared to junior ophthalmologists and even non-specialist opticians. In comparison, RetinaVLM-Specialist’s reports reduced the number of incorrect urgent referrals by almost four times compared to opticians and had higher recall for urgent referrals than junior ophthalmologists. Finally, RetinaVLM is able to reinforce its decisions by citing observable biomarkers within the written report, and highlighting their corresponding regions within the image.
We postulate that the poor performance of existing medical foundation models stems from their lack of detailed knowledge related retinal OCT and AMD. Current VLMs are trained on broad, unstructured datasets that are extracted from medical textbooks, scientific publications or social media posts of healthcare professional [9]. In the United Kingdom and the United States, clinical trainees aspiring to become specialists must undergo up to ten years of post-graduate training to obtain the grade of a board-certified consultant. We argue that training data of current foundation medical VLMs lacks this specialist knowledge and experience, hindering their effective application to real-world clinical tasks.
A core innovation of our work was the creation of a dedicated training curriculum that specializes VLMs in image-based clinical decision making. Analogously to current medical education, this curriculum deconstructs clinical problems into sets of mandatory capabilities required for their resolution and selectively trains VLMs in these skills. To this end, we obtained a large number of tabular reports by processing of retrospectively collected clinical data using advanced algorithms. Additionally, we tasked ophthalmologists to produce a limited number of highly specific textual reports. In total, our curriculum comprises 41,926 OCT images with 479,710 corresponding visual questions and answers. While still modest in size compared to substantially larger foundation datasets, we believe such curated needs-driven approaches are required to deploy language models specialist healthcare. In a similar vein, leading technology companies in artificial intelligence have also started to look beyond the internet’s image and text data to source specialized training data to train LLMs and VLMs in disciplines such as computer programming, journalism, mathematics [24, 25].
Limitations and future research directions
Naturally, the quality of our curriculum depends on the underlying reports, in particular that of the 330 collected detailed textual reports. The majority of the reports used to create RetinaVLM-Specialist were written by a junior ophthalmologist with three years experience. The remaining reports were written by an ophthalmologist with ten years of experience, and their reports were more comprehensive. To match or surpass the performance of the average junior ophthalmologist, we seek to finetune this model further on a small number of reports written by intermediate and senior ophthalmologists.
Additionally, the data may reflect local clinical definitions and workflows. While RetinaVLM-Specialist was trained on reports written by the two UK-based ophthalmologists, it was tested on labels derived from four separate ophthalmologists from different countries and hospitals. In our testing we identified differences in the staging definitions used by the UK-based and Austria-based ophthalmologists, especially with regard to identifying fibrovascular features that differentiate intermediate from inactive late wet AMD. This discrepancy resulted in a number of the misclassifications by RetinaVLM.
Similarly, we observed a discrepancy in image interpretation between junior and senior ophthalmologists. Junior ophthalmologists did not recommend patients for referral if it was likely that the retinal fluid observed was caused by traction rather than neovascularization, as it is not treatable with antiangiogenic drugs. Conversely, the senior ophthalmologist preferred that these patients be still referred for immediate assessment to rule out neovascularization. RetinaVLM was explicitly instructed to refer patients with any sign of fluid of any cause (see Section 5.8.3), and correctly referred more patients as a result.
We found that the LLM generating the question-answer pairs from the reports was sensitive to the specifics of the generation instruction. Extensive trial and error were required to arrive at several instructions, listed in Section A.2, that resulted in diverse sets of high quality question-answer pairs. We discern that all aspects of dataset creation - deciding on the required capabilities, collecting specialized annotations and converting these to question-answer pairs - should be formalized to systematically compare different approaches and ultimately scale dataset curation in the future.
Beyond the formalization and extension of the creation of the curriculum, there are other potential technical improvements to RetinaVLM. Currently, RetinaVLM processes a single two-dimensional OCT image from one type of OCT scanner. In ophthalmological practice, decisions are made based on three-dimensional images from multiple time points, although many recent studies on the use of foundation models in ophthalmology also analyze two-dimensional images [18]. We mitigated the impact of this discrepancy on our study by tasking ophthalmologists to select the most relevant two-dimensional slice of the imaged volume before proceeding with the referral decision. In the future, a more sophisticated vision encoder, which is able to handle three-dimensional data from diverse OCT imaging devices, could be integrated with RetinaVLM. Similarly, one may opt to incorporate multimodal information, such as health questionnaires, clinical tests or the patient’s medical history, into the decision making process [26]. The fundamental model architecture and training would remain similar, but the level of reasoning required for differential diagnosis across multiple scans would potentially increase.
RetinaVLM also inherits some of the fundamental limitations of language models. LLMs are prone to confidently present false or fabricated information, termed hallucinations, which has been identified as problematic in medical contexts [13, 14]. RetinaVLM occasionally hallucinates observations of retinal fluid and consequently diagnoses more advanced AMD stages than necessary. RetinaVLM’s output was also sensitive to the wording of questions and instructions. While this had little impact on our qualitative analysis, extensive trial and error was necessary to ensure that RetinaVLM responded with one of the provided options in the quantitative analyses.
In this study, we exclusively trained RetinaVLM for the management of AMD from OCT images, ignoring other retinal pathologies, such as diabetic retinopathy or glaucoma, or imaging modalities, such as color fundus photography [27, 28]. While this enabled us to explore the potential to encode advanced clinical levels of specialism into VLMs at depth, it would limit the applicability of the current version of our models for ocular screening. We hypothesize that VLMs could be specialized on a wider range of ophthalmological tasks by incorporating additional VQA datasets and training VLMs on them. This requires costly and time-consuming curation of specialized training datasets by medical experts. However, we believe that the involvement of medical experts is necessary as routine clinical skills and patient management protocols are rarely documented in existing datasets used to train AI models.
Conclusion
Foundation vision-language models (VLMs) have the potential to revolutionize healthcare by automatically interpreting medical images and communicating their findings in detailed written reports. Trained on large datasets containing millions of medical images and textual annotations, foundation language models have stirred considerable interest for their expert-level performance on medical licensing exams and case studies. However, in this work we have shown that foundation medical VLMs substantially underperform human experts on routine clinical tasks.
We hypothesize that the training data of foundation medical VLMs currently lacks specialist clinical knowledge and experience. To address this, we developed a curriculum-based approach that integrates the expertise of domain specialists into the training of medical VLMs. The resulting model, RetinaVLM, can produce detailed imaging reports that make accurate recommendations for the clinical management of AMD. It approaches and often matches the performance of junior ophthalmologists in disease staging, and outperforms non-specialist opticians in patient referral.
These results indicate that merely increasing the scale of training datasets is insufficient for the development of VLMs with real-world clinical utility. Instead, medical VLMs require high-quality data directly related to the challenges faced by clinicians in their daily practice. We believe our proposed curriculum-based approach provides a blueprint for specializing VLMs that generate true value in healthcare.
A Supplementary material
A.1 Question-answer generation prompts (introduction to retina)
The limited scope of the first part of the tabular reports led us to use only one prompt (see Section A.1.1) to create the 408,545 question-answers in curriculum part 1, ‘introduction to retina’.
A.1.1 General question-answering module
I am constructing a dataset to train a model to answer questions based solely on OCT images.
The model will access ONLY the image to deduce attributes. For context, the image in question has attributes as follows:
<ReportText>
Based on these attributes, generate a numbered list of diverse questions and answers.
Ensure the format is:
1. Q: [Question about an image attribute]
A: [Specific answer deduced from the image]
Rules:
- Questions should be crafted in a way that they don’t explicitly state the attribute values, but the answers should be based on them. All attributes can be determined from the image.
- Incorporate both yes/no and open-ended styled questions, but always provide a definitive answer in the answer section.
- Occasionally touch on patient outcome/treatment.
A.2 Question-answer generation prompts (advanced retinal specialism)
Creating a large quantity and variety of question-answer pairs was important for preventing overfit when finetuning on the 330 images that constitute curriculum part 2. The prompts each focus on a different aspect of question answering and report writing, but are expected to have some overlap in the question-answer pairs they create. The six modules used to generate up to 230 question-answers per image report, as shown in Figure 2, are covered by the following 10 prompts:
  1. 1.
    Advanced biomarkers (Up to 30 total question-answers)
    • Up to 30 question-answer pairs specific to biomarkers (see Section A.2.1)
  2. 2.
    Disease staging definitions (Up to 50 total question-answers)
    • Up to 20 question-answers that use the observation and staging guidelines (see Section A.2.2)
    • Up to 30 question-answers that train the model to replicate any reasoning linking observations to the disease stage that is present in the report (see Section A.2.3)
  3. 3.
    Staging reasoning (Up to 60 total question-answers)
    • Up to 20 question-answers that use the observation and staging guidelines to train the model to explain the diagnosed disease stage in terms of the visible biomarkers (see Section A.2.4)
    • Up to 40 question-answers that focus on training the model in differentiating active from inactive late wet AMD (see Section A.2.5)
  4. 4.
    Referral reasoning (Up to 25 total question-answers)
    • Up to 25 question-answer pairs that focus on training the model to reason about different levels of urgency in patient referral (see Section A.2.6)
  5. 5.
    General visual question-answering (Up to 30 total question-answers)
    • Up to 15 question-answers created directly from the report, but specific to one attribute (see Section A.2.7)
    • Up to 15 question-answers created directly from the report, but including longer and more open-ended questions A.2.8)
  6. 6.
    Report writing (Up to 35 total question-answers)
    • Up to 15 question-answers that focus on replicating the specialist’s report (see Section A.2.9)
    • Up to 20 question-answers that all three observational, disease staging and patient referral guidelines to train the model to create more sophisticated reports A.2.10)
A.2.1 Advanced biomarkers module
  I am constructing a dataset to train a model to answer questions based solely on OCT images.  Below are guidelines outlining what can appear in an OCT image:  <ObservationGuidelines>  However, the image the model is being asked about is characterised by the following description:  DESCRIPTION OF IMAGE: "<ReportText>"  Task: Write 30 varied questions and answers that ask the model about the image.  Ensure the format is:    1. Q: [Question or statement to describe the image]    A: [Augmented version of actual image description]  Rules:  - Ask separately about the presence, amount, location and type of some of the biomarkers in the guidelines  - Ask about the presence or absence of the biomarkers in the guidelines  - Rather than saying the image/description does not specify/mention a biomarker, instead say ’the image does not show/display/exhibit evidence of’ the biomarker UNLESS its presence is already implied by another present biomarker. The model does not see the above description of the image, it’s only given the original image when answering questions.  - Try not to give away too much information included in the image description in the question text. The model must learn to use the image to determine the answer, and not make educated guesses based on the question alone.  - The answer must be accurate and reflect the same information in the image description.  - Write nothing except the questions and answers.  Tips:  - Example questions: "Does this image show any subretinal fluid?" "Do you see any intraretinal fluid?" "Is there any hypertransmission? "Does the image show a PED?"  - Answer style variation: Finally, do not start too many questions with "No, ..." or "Yes, ...". Vary the answer style (the biomarker is ’not present’, ’is shown’, ’exhibits no’, ’does contain’ etc...)  - Positive and negative balance: Try to include an even balance of questions with positive responses (i.e. ask the model about each of the biomarkers that were reported in the description) and negative responses (i.e. that biomarker is not present)    To do this, if you create a question about a biomarker that isn’t in the description, try to create a second, similar question about a biomarker that is observable in the image.    In order to make the question set not give too much away about the image, you can make paired questions which have positive and negative responses.    For example, for an image with a PED but no subretinal fluid, if you ask f.e.      "Q: Is there a PED? If so, where? A: There is a PED present, it’s in the center..."    you should also create a question with a negated answer f.e.      "Q: Is there any subretinal fluid? And in what quantity? A: There is no sign of subretinal fluid in the image...".    This will help you keep an even balance of positive and negative responses, so that the model cannot guess the answer to the question without considering the image.
A.2.2 Disease staging definitions modules
  I am constructing a dataset to train a model to answer questions based solely on OCT images.  Below are guidelines outlining what can appear in an OCT image:  <ObservationGuidelines>  <DiseaseStagingGuidelines>  However, the image the model is being asked about is characterised by the following description:  DESCRIPTION OF IMAGE: "<ReportText>"  Task: Write 20 varied questions and answers that ask the model about the image.  Ensure the format is:    1. Q: [Question or statement to describe the image]    A: [Augmented version of actual image description]  Rules:  - Information about the image should not be in the question.  - The answer must be accurate and reflect the same information in the image description.  - Write nothing except the questions and answers.  Tips:  - Questions should be specific and ask about certain attributes, or sets of attributes. For example "Q: Is there any subretinal fluid in this image?" or "Q: Is the AMD stage intermediate, or is it more advanced?".  - Ask separately about the presence, amount, location and type of some of the biomarkers. Try to create an even balance of ’yes and ’no’ answers.  - Rather than saying the image does not directly/explicitly specify/mention the presence of an attribute, instead say ’the image does not exhibit/show/display/evidence’ the attribute, UNLESS its presence is already directly implied by the presence of another attribute. The model does not see the above description of the image, it’s only given the original image when answering questions.  - Sometimes the desired output format should be specific in the question (f.e. Answer with ’yes’ or ’no’, or answer by stating if the image ’does’ or ’does not’ contain the biomarker in question.)
A.2.3 Disease staging from the report module
  I am constructing a dataset to train a model to answer questions based solely on OCT images.  Common disease stages for age-related macular degeneration are: healthy (no-AMD), non-AMD pathology, early AMD, intermediate AMD, late dry AMD, late wet (inactive) AMD, late wet (active) AMD  These stages may be referred to with slightly different names in the image description.  If the description doesn’t specify whether the late wet AMD is inactive or active, then you shouldn’t either (just see how to description refers to it, but don’t feel the need to copy it verbatim)  However, the image the model is being asked about is characterised by the following description:  DESCRIPTION OF IMAGE: "<ReportText>"  Task: Write 30 varied questions and answers that require the model to estimate the patient’s disease stage from the image.  Ensure the format is enumerated:    1. Q: [Question or statement to describe the image]    A: [Augmented version of actual image description]  Rules:  - If the estimated disease stage is not provided in the report, do not write ANY questions and answers. Simply write "No disease stage in report".  - Try not to give away too much information included in the image description in the question text. The model must learn to use the image to determine the answer, and not make educated guesses based on the question alone.  - The answer must be accurate and reflect the same information in the image description.  - The questions and answers must vary in their style, formulation and vocabulary.  - Write nothing except the questions and answers.  Tips:  - Most questions should ask the model to first estimate the disease stage    For example, "Decide the most advanced AMD stage supported by the image, and explain your reasoning by noting any biomarkers most relevant to that stage."  - If the explanation or reasoning for the disease stage is given in the report, make sure to include that in the model’s answers  - Questions 1 to 20 should first ask the model to decide/estimate/determine/identify/... the (most advanced) disease stage, and then explain their answer.  - Questions 21 to 30 questions should ask for the biomarkers and then the disease stage, such as "Describe any relevant/notable/significant biomarkers, and link them to the most likely disease stage" (which will be the one in the report)
A.2.4 Disease staging reasoning (with guidelines)
  I am constructing a dataset to train a model to answer questions based solely on OCT images.  Below is a schema outlining what can appear in an OCT image:  <ObservationGuidelines>  <DiseaseStagingGuidelines>  However, the image the model is being asked about is characterised by the following description:  DESCRIPTION OF IMAGE: "<ReportText>"  Task: Write 20 varied questions and answers that require the model to perform chain-of-thought reasoning.  Ensure the format is:    1. Q: [Question or statement to describe the image]    A: [Augmented version of actual image description]  Rules:  - Information about the image listed in the description should not be used in the question, only in the answer.  - Never use the word ’description’ or ’mention’ in the answer, the model does not see the description of the image, it only sees the original image itself.  - The answer must be accurate and reflect the same information in the image description.  - Write nothing except the questions and answers.  Tips:  - Some questions should ask the model to provide a long and detailed answer to questions like ’Describe all the observable biomarkers in the image and then link these to the most likely disease stage’.  - One or two questions should ask the model to explain/deduce the highest precedent AMD stage (i.e. the overall AMD stage) by describing the biomarker(s) in the image which belong to the most advanced disease stage, and linking them to the relevant disease stage.  - Some questions should ask the model to summarise any relevant biomarkers and, explaining its reasoning using the guidelines, conclude with the AMD stage. For example, a drusenoid PED suggests intermediate AMD, but if coupled with subretinal fluid the overall AMD stage becomes active late wet due to the fluid.  - Some questions should ask the model to fully describe and list all its image observations, and then conclude the presence, absence, location or quantity of a specific biomarker (f.e. ’Describe the OCT image in detail and note any abnormalities, and then tell me if the image contains subretinal fluid.’)
A.2.5 Disease staging reasoning second module
   I am constructing a dataset to train a model to answer questions based solely on OCT images.  Below are guidelines outlining what can appear in an OCT image:  <DiseaseStagingGuidelines>  However, the image the model is being asked about is characterised by the following description:  DESCRIPTION OF IMAGE: "<ReportText>"  Task: Write 40 varied questions and answers that require the model to estimate the patient’s disease stage from the image.  Ensure the format is enumerated:    1. Q: [Question or statement to describe the image]    A: [Augmented version of actual image description]  Rules that always apply:  - If the estimated disease stage is not provided in the report, do not write ANY questions and answers. Simply write "No disease stage in report".  - If reasoning for the disease stage is provided in the report, you MUST include this logic/nuance in the model’s answers.  - You MUST not give away information about the image description in the question text. The model must learn to use the image to determine the answer, and not make educated guesses based on the question alone.  - The answer must be accurate and reflect the same information in the image description.  - The questions and answers must vary in their style, formulation and vocabulary.  - Write nothing except the questions and answers.  Rules that are specific to differentiating cases of active, vs inactive, late wet AMD:  - In cases with an active late wet diagnosis, you MUST make it clear that the PRESENCE OF FLUID is what differentiates active from inactive late wet AMD. Make this clear when explaining the reasoning behind active late wet diagnoses.    So do not imply that f.e. "The disease stage is late wet AMD (active), as indicated by the subretinal/intraretinal fluid, subretinal hyperreflective material, ...."    Instead, you MUST explain that f.e. "The overall disease stage is late wet AMD, which is active due to the detection/presence of subretinal/intraretinal fluid. Inactive late wet features include fibrovascular PED, ..."    Or "Late wet AMD best describes the AMD stage. The detection/presence of subretinal/intraretinal fluid means this is active late wet AMD. Other late wet features include fibrovascular PED, ..."  - Similarly, in cases with an inactive late wet diagnosis, you MUST make it clear in the model’s reasoning that the lack of fluid, in combination with the other biomarkers, is what resulted in the inactive late wet diagnosis.    For example, "The stage is late wet AMD according to the evidence/presence of subretinal hyperreflective material, but it is inactive as there is no detectable fluid of any kind in the image"  - The exact formulation of this answer must vary according to the question. Do not copy the examples too many times. Add a lot of diversity in the model’s responses.  Tips:  - Most questions should ask the model to first estimate the disease stage  - Some questions should first ask the model for its disease stage prediction, and then ask it to explain its reasoning by noting visible biomarkers that relate to that stage    For example, "Decide the most advanced AMD stage supported by the image, and explain your reasoning by noting any biomarkers most relevant to that stage."  - The explanation for the disease stage may already be given in the report, but you can also use the guidelines provided to work out which biomarkers resulted in that disease stage.  - Questions 31 to 40 questions should ask for the biomarkers and then the disease stage, such as "Describe any relevant/notable/significant biomarkers, and link them to the most likely disease stage" (which will be the one in the report)
A.2.6 Patient referral reasoning module
  I am constructing a dataset to train a model to answer questions based solely on OCT images.  Below are guidelines outlining what can appear in an OCT image:  <DiseaseStagingGuidelines>  <PatientReferralGuidelines>  However, the image the model is being asked about is characterised by the following description:  DESCRIPTION OF IMAGE: "<ReportText>"  Task: Write 25 varied questions and answers that require the model to perform explain its reasoning before making conclusions and recommendations.  Ensure the format is:    1. Q: [Question or statement to describe the image]    A: [Augmented version of actual image description]  Rules:  - Information about the image should not be in the question.  - Never use the word ’description’ or ’mention’ in the answer, the model does not see the description of the image, it only sees the original image itself.  - The answer must be accurate and reflect the same information in the image description.  - Write nothing except the questions and answers.  - The model sees each questions separately so they will not be seen together.  Tips:  - Some questions should ask the model to list any relevant biomarkers and, based on these, recommend if the patient should be referred or not.  - Many questions should make a series of requests, by asking the model to write a long and detailed answer reporting all the observable biomarkers, linking those to the most likely disease stage and then summarising the report with a patient referral recommendation.    For example, ’Describe the all biomarkers in the image, and based off of your observations which AMD best describe the patient. Summarise your report with a referral recommendation that follows the treamtent guidelines.’  - Some questions should ask the model to recommend if the patient does: not need referral, if they need general attention by a specialist, or if they likely need treatment with anti-vegf based off the models observations  - Some questions should ask the model to explain/deduce/estimate the patient’s risk (i.e. the recommended referral action) by first describing the most advanced or concerning biomarker(s) (i.e. the observable biomarker(s) which belong to the most severe disease stage)
A.2.7 Specific report-based questions module
  I am constructing a dataset to train a model to answer questions based solely on OCT images.  The image in question is characterised by the following description:  DESCRIPTION OF IMAGE: "<ReportText>"  Task: Write 15 varied questions and answers that ask the model about the image. They should  Ensure the format is:    1. Q: [Question or statement to describe the image]    A: [Augmented version of actual image description]  Rules:  - Try not to give away too much information included in the image description in the question text. The model must learn to use the image to determine the answer, and not make educated guesses based on the question alone.  - The answer must be accurate and reflect the same information in the image description.  - Write nothing except the questions and answers.  Tips:  - Questions should be specific and ask about certain attributes, or sets of attributes. For example "Q: Is there any subretinal fluid in this image?".
A.2.8 General report-based questions module
    I am constructing a dataset to train a model to answer questions based solely on OCT images.  The image in question is characterised by the following description:  DESCRIPTION OF IMAGE: "<ReportText>"  Task: Write 15 varied questions and answers that ask the model about the image.  Ensure the format is:    1. Q: [Question or statement to describe the image]    A: [Augmented version of actual image description]  Rules:  - The answer should be a modified and augmented version of the actual description.  - Try not to give away too much information included in the image description in the question text. The model must learn to use the image to determine the answer, and not make educated guesses based on the question alone.  - The answer must be accurate and contain the same information as the actual description. However, the order of the sentences as they are written must change and randomly vary.  - Write nothing except the questions and answers.  Tips:  - Some question should be general and ask to describe the image.  - Other questions should be more specific such as "Q: Is there any subretinal fluid in this image?" that have shorter answers.  - Example questions/statements might be "Describe the OCT image in detail." or "Can you give me a summary of the image?".
A.2.9 Basic report writing module
  I am constructing a dataset to train a model to answer questions based solely on OCT images.  The image in question is characterised by the following description.  DESCRIPTION OF IMAGE: "<ReportText>"  Task: Write 15 questions and answers that ask the model to describe the image in full. The first ten questions should ask to describe the entire image (with jumbled/permuted/randomised answers), while the final five should be more specific or use segments of the description.  Ensure the format is:    1. Q: [Question or statement to describe the image]    A: [Augmented version of actual image description]  Rules:  - The answer should be a modified and augmented version of the actual description.  - The answer should contain the same information as the actual description but the order of the sentences as they are written must change and randomly vary  - The question must not contain any information about the image.  - Write nothing except the questions and answers.  Tips:  - Example questions/statements might be "Describe the OCT image in detail." or "Can you give me a summary of the image?" or "Write a report on this image to be given to an optometrist.".
A.2.10 Advanced report writing module
  I am constructing a dataset to train a model to answer questions based solely on OCT images.  Below are guidelines outlining what can appear in an OCT image:  <ObservationGuidelines>  <DiseaseStagingGuidelines>  <PatientReferralGuidelines>  However, the image the model is being asked about is characterised by the following description:  DESCRIPTION OF IMAGE: "<ReportText>"  Task: Write 20 requests for reports that ask the model reason all the way from the observable biomarkers, linking these to the most advanced disease stage, and finally to a referral recommendation for the patient.  Ensure the format is:    1. Q: [Request to the model]    A: [Answer or report deduced using actual image description]  Rules:  - Never use the word ’description’ or ’mention’ in the answer, the model does not see the description of the image, it only sees the original image itself.  - Try not to give away too much information included in the image description in the question text. The model must learn to use the image to determine the answer, and not make educated guesses based on the question alone.  - Write nothing except the questions and answers.  - The model sees each questions separately so they will not be seen together.  Tips:  - Example questions/statements might be "Describe the OCT image in detail." or "Can you give me a summary of the image?" or "Write a report on this image to be given to a retinal specialist.".  - Ask the model to explain why the AMD stage is not more advanced than it is (i.e. the absence of certain late stage biomarkers)  - Ask the model to explain why the AMD stage is more advanced than an earlier stage (i.e. the presence of certain late stage biomarkers)  - A few times, request verbose reports like "Write a report that starts by highlighting the most significant and salient biomarkers, and link these to the most probable disease stage. Conclude with a referral recommendation."